OpenStudy (anonymous):

During the first part of a canoe trip, Ken covered 60km at a certain speed. If he then traveled at 24km at a speed that was 4km/h slower . If the total time for the trip was 8 hours, what is the speed on each part of the trip.?

OpenStudy (anonymous):

Since you have two unknowns, you need two equations to solve this problem. The first equation is fairly simple. It's just: x+y=8 Where x= the time for the first part of the trip and y= time for the second part of the trip, The second equation is a little bit tricky. So let 60/x = speed of the first part. Let 24/y = the speed of the second part of the trip. Since the speed of the first part of the trip is equal to the speed of the second part of trip + 4 Km/hr, we can write the second equation is follows: 60/x = 24/y + 4 Now we have two equations; so we can now solve. Rewrite the first equation to look like this: x=8-y Now substitute into the other equation. 60/(8-y) = 24/y +4 Multiply both sides by (8-y) 60 = (24/y +4) (8-y) 60 = (192-24y)/y +32 -4y 28 = (192 -24y)/y -4y Multiply both sides by y 28y = 192 -24y -4y^2 4y^2 +52y -192 Solve for y using quadratic equation\[(-52\pm \sqrt{52^{2}-(4*4*-192)})/(2*4)\] So y = 3 hours So x = 8-3 or 5 hours So the speed for the first part of the trip is 60km/5hr or 12km/hr The speed for the second part of the trip is 24km/3hr or 8 km/hr